# Source-Free Domain Adaptation with Diffusion-Guided Source Data Generation

### v1

@@ 4.7.1 Bridging Domain Spaces (FixBi)
| ResNet-101 (He et al., 2015) | ✗ | 55.1 | 53.3 | 61.9 | 59.1 | 80.6 | 17.9 | 79.7 | 31.2 | 81.0 | 26.5 | 73.5 | 8.5 | 52.4 |
| MCC (Jin et al., 2019) | ✗ | 88.7 | 80.3 | 80.5 | 71.5 | 90.1 | 93.2 | 85.0 | 71.6 | 89.4 | 73.8 | 85.0 | 36.9 | 78.8 |
| GSDA (Hu et al., 2020) | ✗ | 93.1 | 67.8 | 83.1 | 83.4 | 94.7 | 93.4 | 93.4 | 79.5 | 93.0 | 88.8 | 83.4 | 36.7 | 81.5 |
| FixBi (Na et al., 2021) | ✗ | 96.1 | 87.8 | 90.5 | 90.3 | 96.8 | 95.3 | 92.8 | 88.7 | 97.2 | 94.2 | 90.9 | 25.7 | 87.2 |
| CoVi (Na et al., 2022) | ✗ | 96.8 | 85.6 | 88.9 | 88.6 | 97.8 | 93.4 | 91.9 | 87.6 | 96.0 | 93.8 | 93.6 | 48.1 | 88.5 |
| SHOT (Liang et al., 2020) | ✓ | 94.3 | 88.5 | 80.1 | 57.3 | 93.1 | 94.9 | 80.7 | 80.3 | 91.5 | 89.1 | 86.3 | 58.2 | 82.9 |
| 3C-GAN (Li et al., 2020) | ✓ | 94.8 | 73.4 | 68.8 | 74.8 | 93.1 | 95.4 | 88.6 | 84.7 | 89.1 | 84.7 | 83.5 | 48.1 | 81.6 |
| NRC (Yang et al., 2021) | ✓ | 96.8 | 91.3 | 82.4 | 62.4 | 96.2 | 95.9 | 86.1 | 80.6 | 94.8 | 94.1 | 90.4 | 59.7 | 85.9 |
| NRC++ (Yang et al., 2023) | ✓ | 96.8 | 91.9 | 88.2 | 82.8 | 97.1 | 96.2 | 90.0 | 81.1 | 95.2 | 93.8 | 91.1 | 49.6 | 87.8 |
| DM-SFDA (FixBi) | ✓ | 95.9 | 87.6 | 89.0 | 86.2 | 96.6 | 95.0 | 91.3 | 86.4 | 96.5 | 94.0 | 90.4 | 30.2 | 86.6 |
| DM-SFDA (CoVi) | ✓ | 96.3 | 86.0 | 88.7 | 85.4 | 97.1 | 93.4 | 91.0 | 86.0 | 95.8 | 93.7 | 91.3 | 46.6 | 87.5 |
4.7.2 Contrastive Viscinal Space (CoVi)
CoVi (Na et al., 2022) utilises the viscinal space between the source and target domain for unsupervised domain adaptation. A Contrastive Viscinal Space-based algorithm is introduced, where the viscinal space is divided into contrastive space (source dominant) and consensus space (target dominant) by using the Entropy Maximization Point (EMP) as a boundary to resolve the problem of equilibrium collapse of labels. In the contrastive space, the inter-domain discrepancy is mitigated by constraining instances to have contrastive views and labels. The consensus space reduces the confusion between intra-domain categories by enforcing label consensus among the multiple target-dominant vicinal instances to a single target label.
@@ 5.1 Datasets
Office-31: Office-31 (Saenko et al., 2010) is a benchmark image classification dataset that consists of a limited set of images distributed across 31 categories spanning three domains: Amazon (2,817 images), DSLR (498 images), and Webcam (795 images).
Office-Home: Office-Home (Venkateswara et al., 2017), on the other hand, comprises a more extensive dataset with a total of 15.5K images from 65 classes, gathered from 4 distinct image domains: Artistic, Clipart, Product, and Real-world. Our analysis includes 12 transfer tasks for this dataset.
VisDA: VisDA (Peng et al., 2017) encompasses two distinct domains: synthetic and real, each comprising 12 classes. The synthetic domain holds around 150K computer-generated 3D images with various poses, while the corresponding real domain includes approximately 55K images captured from the real world.
5.2 Evaluation Metric
Our evaluation metrics encompass per-class accuracy and their average (Avg.) performance. Our evaluation involves transfer tasks spanning all the domains of each dataset, and we report the top-1 accuracy achieved by various methods under different settings.
@@ 5.3 Experimental Setup
We implement our approach using PyTorch and use ResNet-50 (He et al., 2015) as the backbone network for the Office-31 (Saenko et al., 2010) and Office-Home (Venkateswara et al., 2017) datasets and Resnet-101 for the VisDA (Peng et al., 2017) dataset. All the experiments for our proposed approach were conducted on a Nvidia A100 GPU. The other details for specific parts of our pipeline are specified in the subsequent sub-sections.
5.3.1 Diffusion Model Fine-tuning
We use Stable Diffusion v1.4 (Rombach et al., 2022) as the base model for all experiments. The finetuning of the diffusion models was done in the Accelerate (Gugger et al., 2022) environment. Memory-efficient attention was enabled using xFormers (Lefaudeux et al., 2022) for all the experiments performed. Furthermore, we used the Low-Rank Adaptation (LORA) of the Stable Diffusion Pipeline.
5.3.2 Denoising Diffusion Policy Optimization
In the context of the DDPO experiments, we employed the Low-Rank Adaptation (LORA) technique within the framework of the Stable Diffusion Pipeline. The training was done for a batch size of 4, and 100 batches were sampled per step. The training procedure encompassed 100 steps, with each step incorporating two distinct phases: a sampling phase and 10 consecutive inner training epochs dedicated to training the model on the sampled data from the previous phase.
5.3.3 Unupervised Domain Adaptation
In the unsupervised domain adaptation phase of our pipeline, we use the experimental setup proposed in each individual approaches (Na et al., 2021, 2022). For the UDA approaches, we use the generated source data labeled using the pre-trained model and all unlabeled target data. We use minibatch stochastic gradient descent (SGD) with a momentum of 0.9, an initial learning rate of 0.001, and a weight decay of 0.005. We follow the same learning rate schedule as in (Ganin & Lempitsky, 2015a).
@@ 5.4.1 Results on Office-31
We present the outcomes for the Office-31 dataset in Table 1. By employing the source data produced through our proposed pipeline as input, the UDA methods managed to surpass the existing state-of-the-art SFDA approaches by margins of 0.6% and 0.7%, achieving average accuracies of 90.1% and 90.2% with the FixBi (Na et al., 2021) and CoVi (Na et al., 2022) approaches respectively.
@@ 5.4.2 Results on Office-Home
The results for the Office-Home dataset (Venkateswara et al., 2017) are presented in Table 2. By integrating the source data generated through our proposed pipeline into the UDA methodologies, we have successfully surpassed existing methods, achieving an average accuracy of 72.4% and 72.6% with FixBi (Na et al., 2021) and CoVi (Na et al., 2022) respectively.
@@ 5.4.3 Results on VisDA17
We summarize the results for the VisDA dataset (Peng et al., 2017) in Table 3. Our proposed framework is able to outperform the source-free baselines across many classes like bus and car by 0.8% and 3.4% respectively. The accuracies of our proposed method are very close to the approaches where source data is available, thereby showing the effectiveness of our data generation pipeline.
@@ 6.1 Selective Pseudo Labeling Target Data
| ✗ | 67.8 | 68.3 | 60.1 | 95.4 | 60.5 | 98.7 | 75.13 | ||
| ✓ | 93.6 | 94.0 | 77.0 | 99.0 | 78.0 | 99.9 | 90.2 |
6.2 Image Caption Generators
The output of a text-to-image diffusion model is conditioned on the input prompt. To derive the optimal output from the diffusion model, we used an implicit captioner as described in Section 4.4. In this context, we explored the use of an off-the-shelf Vision Language Model, LlaVa (Liu et al., 2023), to produce captions from images, which then served as inputs for the diffusion models. The outcomes from the caption generator ablation study are consolidated in Table 5. Our findings indicate that fine-tuning an implicit captioner yields superior results compared to utilizing an off-the-shelf model.
@@ 6.2 Image Caption Generators
| LLaVa | 90.5 | 92.0 | 75.2 | 95.4 | 76.1 | 99.5 | 88.1 | |
| Implicit Captioner | 93.6 | 94.0 | 77.0 | 99.0 | 78.0 | 99.9 | 90.2 |
7 Discussion
7.1 Visualization Analysis
Figure LABEL:fig:tsne illustrates the t-SNE visualizations (van der Maaten & Hinton, 2008) for the latent feature space of the target domain, post-adaptation, using the Office-31 dataset (Saenko et al., 2010). It is evident that the target instances are significantly clustered alongside the source instances within each class following the adaptation process. This clustering underscores the model’s capability to perform accurate class conditional generation within the target domain, thereby showcasing its superior performance in data generation. To further demonstrate the effectiveness of our proposed approach, we visually inspect the generated source domain images. Figure 3 shows the data generated for the Office-31 dataset for the DSLR Amazon task. As shown, the proposed approach was able to successfully generate images similar to the source domain.
7.2 Challenges and Limitations
@@ 8 Conclusion
9 Impact Statement
The proposed work introduces a novel approach for SFDA through the regeneration of source data using a text-to-image diffusion model. The implications of this work span several dimensions of societal, ethical, and technological realms, promising to influence various sectors positively while also presenting areas requiring careful consideration.
9.1 Societal Impact
Our methodology has the potential to democratize access to advanced machine learning capabilities by reducing the dependency on large, annotated datasets. By enabling effective domain adaptation without direct access to source data, we can foster innovation in areas where data privacy concerns or proprietary restrictions limit data availability. Applications in healthcare, environmental monitoring, and autonomous systems could see accelerated advancements, improving services and outcomes in these critical sectors.
9.2 Ethical Considerations
The use of pre-trained models and generative techniques raises important ethical questions. While our approach aims to mitigate data availability challenges, it also emphasizes the need for vigilance against perpetuating biases present in pre-trained models. Generated images must be scrutinized for biases that could lead to unfair or prejudiced outcomes, especially in sensitive applications. Moreover, the ethical use of generative models must be guided by principles that prevent misuse, such as generating misleading or harmful content.
9.3 Technological Advancements
Beyond immediate applications, this work contributes to the broader field of machine learning by exploring novel uses of diffusion models in domain adaptation. It challenges current paradigms and opens new avenues for research in unsupervised learning, model generalization, and the efficient utilization of pre-trained models. As the technology matures, it could lead to more robust, adaptable systems capable of learning from diverse and limited data sources.

### v2

@@ 4.6 Phase IV: Unsupervised Domain Adaptation
| ResNet-101 [14] | ✗ | 55.1 | 53.3 | 61.9 | 59.1 | 80.6 | 17.9 | 79.7 | 31.2 | 81.0 | 26.5 | 73.5 | 8.5 | 52.4 |
| MCC [19] | ✗ | 88.7 | 80.3 | 80.5 | 71.5 | 90.1 | 93.2 | 85.0 | 71.6 | 89.4 | 73.8 | 85.0 | 36.9 | 78.8 |
| GSDA [16] | ✗ | 93.1 | 67.8 | 83.1 | 83.4 | 94.7 | 93.4 | 93.4 | 79.5 | 93.0 | 88.8 | 83.4 | 36.7 | 81.5 |
| FixBi [33] | ✗ | 96.1 | 87.8 | 90.5 | 90.3 | 96.8 | 95.3 | 92.8 | 88.7 | 97.2 | 94.2 | 90.9 | 25.7 | 87.2 |
| CoVi [34] | ✗ | 96.8 | 85.6 | 88.9 | 88.6 | 97.8 | 93.4 | 91.9 | 87.6 | 96.0 | 93.8 | 93.6 | 48.1 | 88.5 |
| ICON [56] | ✗ | 96.6 | 91.8 | 88.0 | 82.0 | 96.8 | 93.3 | 90.0 | 81.0 | 95.2 | 93.6 | 91.0 | 49.5 | 87.4 |
| SHOT [26] | ✓ | 94.3 | 88.5 | 80.1 | 57.3 | 93.1 | 94.9 | 80.7 | 80.3 | 91.5 | 89.1 | 86.3 | 58.2 | 82.9 |
| 3C-GAN [25] | ✓ | 94.8 | 73.4 | 68.8 | 74.8 | 93.1 | 95.4 | 88.6 | 84.7 | 89.1 | 84.7 | 83.5 | 48.1 | 81.6 |
| NRC [51] | ✓ | 96.8 | 91.3 | 82.4 | 62.4 | 96.2 | 95.9 | 86.1 | 80.6 | 94.8 | 94.1 | 90.4 | 59.7 | 85.9 |
| NRC++ [53] | ✓ | 96.8 | 91.9 | 88.2 | 82.8 | 97.1 | 96.2 | 90.0 | 81.1 | 95.2 | 93.8 | 91.1 | 49.6 | 87.8 |
| AaD [52] | ✓ | 97.4 | 90.5 | 80.8 | 76.2 | 97.3 | 96.1 | 89.8 | 82.9 | 95.5 | 93.0 | 92.0 | 64.7 | 88.0 |
| AaD w/ Co-learn [57] | ✓ | 97.6 | 90.2 | 85.0 | 83.1 | 97.1 | 92.1 | 84.9 | 96.8 | 96.8 | 95.1 | 92.2 | 56.8 | 89.1 |
| DM-SFDA | ✓ | 98.1 | 89.8 | 90.6 | 90.5 | 96.8 | 95.2 | 92.2 | 93.4 | 97.8 | 94.4 | 92.4 | 48.8 | 86.3 |
@@ 5.1 Datasets
Office-31: Office-31 [39] is a benchmark image classification dataset that consists of a limited set of images distributed across 31 categories spanning three domains: Amazon (2,817 images), DSLR (498 images), and Webcam (795 images).
Office-Home: Office-Home [48], on the other hand, comprises a more extensive dataset with a total of 15.5K images from 65 classes, gathered from 4 distinct image domains: Artistic, Clipart, Product, and Real-world. Our analysis includes 12 transfer tasks for this dataset.
VisDA: VisDA [35] encompasses two distinct domains: synthetic and real, each comprising 12 classes. The synthetic domain holds around 150K computer-generated 3D images with various poses, while the corresponding real domain includes approximately 55K images captured from the real world.
@@ 5.2 Experimental Setup
We implement our approach using PyTorch and use ResNet-50 [14] as the backbone network for the Office-31 [39] and Office-Home [48] datasets and Resnet-101 for the VisDA [35] dataset. All the experiments for our proposed approach were conducted on a Nvidia A100 GPU. The other details for specific parts of our pipeline are specified below:
Diffusion Model Fine-tuning: We use Stable Diffusion v1.4 [38] as the base model for all experiments. The finetuning of the diffusion models was done in the Accelerate [12] environment. Memory-efficient attention was enabled using xFormers [23] for all the experiments performed. Furthermore, we used the Low-Rank Adaptation (LORA) of the Stable Diffusion Pipeline.
AlignProp: In the context of the AlignProp experiments, we employed the Low-Rank Adaptation (LORA) technique within the framework of the Stable Diffusion Pipeline. The training was done for a batch size of 4, and 100 batches were sampled per step. The training procedure encompassed 100 steps, with each step incorporating two distinct phases: a sampling phase and 10 consecutive inner training epochs dedicated to training the model on the sampled data from the previous phase.
Unupervised Domain Adaptation: In the unsupervised domain adaptation phase of our pipeline, we use the experimental setup proposed in [33]. For the UDA approach, we use the generated source data labeled using the pre-trained model and all unlabeled target data. We use minibatch stochastic gradient descent (SGD) with a momentum of 0.9, an initial learning rate of 0.001, and a weight decay of 0.005. We follow the same learning rate schedule as in [8].
@@ 5.3.1 Results on Office-31
We present the outcomes for the Office-31 dataset in Table 1. By employing the source data produced through our proposed pipeline as input along with the diffusion-model based mixup strategy, our approach is able to match/outperform existing state-of-the-art SFDA approaches for all the tasks, achieving an average accuracy of 93.7%, which is 0.7% higher then the current state-of-the-art.
@@ 5.3.3 Results on VisDA17
We summarize the results for the VisDA dataset [35] in Table 3. Our proposed framework is able to outperform the source-free baselines across many classes like plane, bus and car by 0.5%, 2.1% and 7.4% respectively. The accuracy of our proposed method surpass the approaches where source data is available, thereby showing the effectiveness of our data generation and domain mixup pipeline.
@@ 6.1 Selective Pseudo Labeling Target Data
| ✗ | 67.8 | 68.3 | 60.1 | 95.4 | 60.5 | 98.7 | 75.1 | ||
| ✓ | 97.7 | 99.0 | 82.7 | 99.3 | 83.5 | 100.0 | 93.7 |
6.2 Unsupervised Domain Adaptation
In order to test the efficacy of our Diffusion-Model based Unsupervised Domain Adaptation approach, we compare the downstream performance of our proposed approach with the existing off-the-shelf UDA approaches. Since we chose a ResNet backbone as the pre-trained source model, we experiment with the current state-of-the-art ResNet-based UDA approaches. As shown in Table 5, our proposed diffusion model-based mixup approach is able to significantly outperform all other UDA approaches. This shows that our mixup approach is able to generate much better intermediate domain to better bridge the huge domain gap.
@@ 6.2 Unsupervised Domain Adaptation
| DM-SFDA (FixBi [34]) | 93.0 | 93.5 | 77.4 | 98.7 | 78.3 | 99.7 | 90.1 | |
| DM-SFDA (CoVi [33]) | 93.6 | 94.0 | 77.0 | 99.0 | 78.0 | 99.9 | 90.2 | |
| DM-SFDA (ICON [56]) | 95.2 | 92.9 | 78.6 | 99.2 | 78.2 | 100.0 | 90.7 | |
| DM-SFDA | 97.7 | 99.0 | 82.7 | 99.3 | 83.5 | 100.0 | 93.7 |
7 Challenges and Limitations